Harrison Company makes two products and uses a traditional costing system in which a single plantwide predetermined overhead rate is computed based on direct labor-hours. Data for the two products for the upcoming year follow:

Rascon Parcel
Direct materials cost per unit $ 13 $ 7
Direct labor cost per unit $ 2.70 $ 4.00
Direct labor-hours per unit .50 .55
Number of units produced 20,000 48,000


These products are customized to some degree for specific customers.    

Required:

1. The company's manufacturing overhead costs for the year are expected to be $1,019,200. Using the company's traditional costing system, compute the unit product costs for the two products. (Round your intermediate calculations and final answers to 2 decimal places.)

  

2. Management is considering an activity-based absorption costing system in which half of the overhead would continue to be allocated on the basis of direct labor-hours and half would be allocated on the basis of engineering design time. This time is expected to be distributed as follows during the upcoming year:

Rascon Parcel Total
Engineering design time (in hours) 4,300 4,300 8,600

Compute the unit product costs for the two products using the proposed activity-based absorption costing system. (Round your intermediate calculations and final answers to 2 decimal places.)
Answer #1

1) Estimated direct labor hour = (20000*.50+48000*.55) = 36400 Hour

Overhead rate = 1019200/36400 = 28 per hour

Rascon Parcel
Direct material 13 7
Direct labor 2.70 4
Overhead 28*.50 = 14 28*.55 = 15.40
Unit product cost 29.70 26.40

2) Calculate unit product cost

Rascon Parcel
Overhead based on labor hour 140000 369600
Overhead based on engineering time 254800 254800
Total overhead 394800 624400
Unit 20000 48000
Overhead cost per unit 19.74 13.01
Direct material 13 7
Direct labor 2.70 4
Unit product cost 35.44 24.01